OpenStudy (anonymous):

-3(2x-1)=27-2x Please show all work

OpenStudy (cwrw238):

first distribute -3 over the parentheses -3*2x + 3 = 27 - 2x -6x + 3 = 27 - 2x now add 2x to both sides of the equation then subtract 3 from both sides finally divide both sides by coefficient of x

OpenStudy (anonymous):

Thank you so much!

OpenStudy (cwrw238):

yw